OpenStudy (anonymous):

Find the roots in the interval -180≤Ɵ≤180 of the following equation a) tan 3/4 Ɵ=0.5 b) sin 2/5 Ɵ°=-0.3

OpenStudy (campbell_st):

so the question is \[\tan(\frac{3}{4} \theta)= 0.5\]

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

find the arc tan \[\frac{3}{4} \theta = \tan^{-1}(0.5)\]

OpenStudy (anonymous):

wow..thanks..i got it